OpenStudy (anonymous):

find the derivative of t^2/√t+1

OpenStudy (abb0t):

quotient rule: low d high minus high d low all over low-low.

OpenStudy (abb0t):

OR you can bring up the denominator since you should know:\[\frac{ 1 }{ a+1 }= (a+1)^{-1}\]

OpenStudy (anonymous):

i used the quotient rule and i believe i am stuck because i stopped at \[\frac{ 2t-\frac{ t ^{2} }{ 2 } }{ (\sqrt{t+1})^{2} }\]

OpenStudy (abb0t):

I think your derivative is wrong for the numerator. you forgot something next to the 2t. \[(\sqrt{t}+1) \times 2t\]

OpenStudy (abb0t):

and also, \[\frac{ 1 }{ \sqrt{t} } = t^{-\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

i simplified that once i written it out.

OpenStudy (anonymous):

\[u=t^2\] \[u"=2t\] \[v=(t+1)^{\frac{ 1 }{ 2 }}\] \[v"=\frac{ 1 }{ 2 }(t+1)^{-\frac{ 1 }{ 2 }}\] \[=\frac{ 1 }{ 2\sqrt{t+1} }\] \[y"=\frac{ (u"\times v)-(u \times v") }{ v^2 }\] \[=\frac{ 2t \sqrt{t+1}-\frac{ t^2 }{ 2\sqrt{t+1} } }{ t+1 }\] \[=\frac{ 4t(t+1)-t^2 }{ 2\sqrt{t+1} }\times \frac{ 1 }{ t+1 }\] \[=\frac{ 4t^2+4t-t^2 }{ 2\sqrt{t+1} }\times \frac{ 1 }{ t+1 }\]

OpenStudy (anonymous):

Continue that.

OpenStudy (anonymous):

Make sure you read what I did.

OpenStudy (anonymous):

And I used the quotient rule.